Can someone help me? Thank you.
I need to make it clear about the hyperbola. X^2-y^2=1 is transformed T(x,y)=(ax+h, by+k). How do I write the equation after transformation?
Answer by Theo(13342) 

start with T(x,y) = x^2 - y^2
when T(x,y) = 1, then the equation becomes x^2 - y^2 = 1
when you want T(ax+h,by+k), then you simply replace x with ax+h and y with by+k and you get:
T(ax+h,by+k) = (ax+h)^2 - (by+k)^2
when T(ax+h,by+k) is equal to 1, then the equation becomes (ax+h)^2 - (by+k)^2 = 1
that's what i think your solution will be.
| |
|